OpenStudy (anonymous):

What is the difference of g and h? g(x) = 9 + x h(x) = 4x

OpenStudy (anonymous):

(g – h)(x) = –9 + 5x (g – h)(x) = –9 + 3x (g – h)(x) = 9 – 3x (g – h)(x) = 9 – 4x

Nnesha (nnesha):

difference mean subtract so subtract

Nnesha (nnesha):

 now solve